# Hybrid Two-Stage Reconstruction of Multiscale Subsurface Flow with Physics-informed Residual Connected Neural Operator

### v1

@@ 4.2 Hierarchical Operator to Capture Global Information
Although the aforementioned model, by embedding physical laws and residual connections into a deep learning framework, can effectively approximate the solutions to partial differential equations (PDEs), its primary convolutional structure, which focuses on local features, limits its ability to capture global interactions and long-range dependencies. To address this limitation, we have integrated a Transformer module. The Transformer enhances the model’s ability to capture global features and long-range dependencies through self-attention mechanisms and positional encoding, thereby improving the accuracy of predicting the pressure field $p_{\theta}(\cdot)$.
@@
[…]
@@
$\text{Multihead}(Q,K,V)=\text{Concat}(head_{1},head2,...,head_{h})W^{O}$ (24)
each $head_{i}$ is computed as $\text{Attention}(Q_{i},K_{i},V_{i})$, processing different linear projection subspaces, and $W^{O}$ is the output projection matrix belonging to $\mathbb{R}^{h\cdot d_{k}\times d}$, where $h$ is the number of attention heads, $d_{k}$ is the dimension of the key and query vectors, and $d$ is the dimension of the output.
@@
$\text{PE}(i,2k)=\sin\left(\frac{i}{10000^{\frac{2k}{d}}}\right),\quad\text{PE}(i,2k+1)=\cos\left(\frac{i}{10000^{\frac{2k}{d}}}\right)$ (25)
@@
In the context of solving Darcy Flow problems, this capability is particularly important. Certain parts of the pressure field in Darcy Flow problems may be influenced by the distribution of permeability and source terms at a distance. Therefore, the architecture of the Transformer model plays a significant positive role in understanding and predicting these complex spatial relationships.
@@ 4.3 Evaluation Metircs for this stage
Since we add physical information to the loss function, it is not appropriate to use the mean square error as an evaluation index directly. To do this, we use the coefficient of determination (R2) and define the average absolute physical residual (AAPR) to evaluate the coefficient of variation of the model and the degree of fit to the data and physical information, which has been demonstrated by Equation.9.
@@
In this section, we propose a novel operator that integrates a convolutional operator with physical residual links and a Transformer operator, designed to predict the pressure field from multi-scale basis functions and permeability fields. We define $\mathcal{G}_{\text{PINN}}$ as the first part of the mapping in Stage 2, mathematically expressed as:
@@
[…]
5 Results
@@
MSE R2
@@
Table 1: Evaluation metrics for different basis functions.
As shown in the data presented in the table, the mean squared error (MSE) of our proposed first-stage model on the test set for different basis functions is 0.0036, 0.0007, 0.0021, and 0.0027, respectively, while the corresponding R2 values are 0.9716, 0.9148, 0.9366, and 0.9542. These values indicate that such performance is acceptable for deep learning models. To further enhance the credibility of our results, we selected samples from the dataset used in this stage, specifically choosing coarse grids that include fractures and those without fractures, and reconstructed the corresponding basis functions for visualization. These results can be observed in the following figures: validation set—Fig.5 and 6; testing set—Fig.7 and 8; training set—Fig.12 and 13. It should be noted, however, that deep learning techniques cannot guarantee perfect learning of features from the data. Therefore, there is a possibility that the selected samples for visualization may include coarse grids with suboptimal reconstruction quality.
Involved Basis Functions Training Validation Testing Total
@@
Table 2: R2 Performance of Different Basis Function Combinations.
For the second stage, reconstruction for the pressure field, the same step to illustrate the results. Tab.2 and 3Tables 2 and 3 present the numerical results of the evaluation metrics used in our study. In this experiment, we combined various basis functions (including Basis 1, which requires no training) with their corresponding permeability fields and input them into the model for training. For our experiments, we randomly selected different numbers and combinations of basis functions for training. From the test set column in Tab.2, it can be observed that the R2 values of our model are all greater than 0.9, indicating that the model performs excellently in fitting the pressure data. Furthermore, as shown in Tab.3, the AAPR values are all at the magnitude of $1\times 10^{-4}$ or smaller, strongly demonstrating that the proposed model not only accurately reconstructs the data but also adheres to the underlying physical laws of real-world phenomena.
Involved Basis Functions Training Validation Testing Total
@@
Table 3: AAPR Performance of Different Basis Function Combinations.
@@
To comprehensively demonstrate the reconstruction capabilities of our method and the output characteristics of several key modules, we combined and presented these components. The corresponding results can be observed in Fig.10, 11 and 14. These visualizations include the permeability fields of high-contrast fractured porous media, the pressure field solutions of the Darcy equation, partial outputs of the preconditioner, and the reconstructed multiscale basis functions on the corresponding grids. This part of the results can fully show some key outputs of our proposed model and reflect the effectiveness of our model to a certain extent.
@@
6 Discussion and Conclusion
In this paper, we proposed a hybrid-model-training two-stage method, developing an innovative neural operator to solve Darcy problem with high-contrast coefficients under the skeleton of mixed GMsFEM. In the first stage, traditional data-driven training method was hired to reconstruct the multiscale basis functions. For the purpose to eliminate the noise and expand the size of input features, we developed a domain transformation-based preconditioner, using Fourier Neural Operator to process the high-contrast permeability fields, together with a hierarchical multiscale feature extractor to capture the information contained. In the second stage, to efficiently solving the pressure field, we used physics-informed training method to achieve our purpose with a small-sample dataset. To utilize the reconstructed basis functions, we concatenate them with permeability fields as the model input. They will be processed by both local (symmetrical residual connections-based accelerator with a V-shaped CNN) and global (Transformer-based for capture spatial information and long-range dependencies) information learner. Both stages achieved good results in the reconstruction of multiscale basis functions and pressure fields.
Compared to existing research, our proposed model demonstrates superior performance. For the reconstruction of basis functions, since our model is primarily trained in a data-driven manner, the evaluation metrics focusing on the data itself are of particular importance. On the test set, our model achieved mean squared error (MSE) values of 0.0036, 0.0007, 0.0021, and 0.0027, and corresponding R2 values of 0.9716, 0.9148, 0.9366, and 0.9542 for different basis functions. In contrast, the results from (Choubineh et al.)[14] reported MSE values of 0.0466, 0.0743, 0.0184, and 0.0154, with corresponding R2 values of 0.8083, 0.6445, 0.8762, and 0.7625. These results indicate that our model achieves better data fitting and exhibits stronger generalization capabilities.
For the reconstruction of pressure fields in the second stage, our model achieved R2 values ranging from 0.9032 to 0.9111 for different basis function combinations. In comparison, (Choubineh et al.)[15] reported R2 values ranging from 0.8456 to 0.9191. While the R2 values are comparable, we place greater emphasis on whether the results align with real-world physical laws. To address this, we incorporated the average absolute percentage residual (AAPR) as an additional evaluation metric. In this context, all combinations exhibited AAPR values on the order of $1\times 10^{-4}$ or smaller. Despite the similar R2values, our model achieves a higher lower bound for R2, indicating that it delivers comparable or better reconstruction performance while better adhering to physical laws.
While the proposed framework performs well in reconstructing multiscale basis functions and pressure fields, several limitations should be noted. The reliance on large, high-quality datasets in the first stage may limit its applicability in scenarios where data is scarce or expensive to acquire. In addition, the inclusion of physical constraints in the second stage, while beneficial for maintaining physical consistency, adds complexity to the training process and requires careful tuning of hyperparameters. The computational overhead introduced by the Transformer block also presents challenges when scaling to larger datasets or higher dimensional domains.
@@ 6 Discussion and Conclusion
Funding
This research did not receive a specific grant from funding agencies in the public, commercial, or non-profit sectors.
Acknowledgment
Thanks to Dr. Jie Chen for his support, resources, and guidance during this study.
@@ Acknowledgment
Appendix A Knowledge of Mixed Generalized Multiscale Finite Element Methods for Darcy’s Equation[11]
Consider the following Darcy’s Flow problem with high-contrast coefficients and nonhomogeneous boundary condition:
$\begin{cases}k^{-1}u+\nabla p&=0,\quad\text{in}\,\Omega\\ \nabla\cdot u&=f,\quad\text{in}\,\Omega\\ u\cdot n&=g,\quad\text{on}\,\partial\Omega\end{cases}$ (32)
where $k$ is high-contrast heterogeneous permeability field, $u$ and $p$ are the Darcy velocity and pressure, respectively. $f$ is the known source term, $g$ is given normal component of Darcy velocity on the boundary, $\Omega=[0,1]^{2}$ is the computational domain, and $n$ is the outward unit norm vector on the boundary.
Let $\mathcal{T}^{H}$ denote a conforming partition of $\Omega$ into finite elements with coarse-grid size $H$ and $\mathcal{T}^{h}$ denote the fine-grid partition of $\Omega$ into non-overlapping elements with size $h$. Define $\mathcal{E}^{H}:=\bigcup_{i=1}^{N_{e}}E_{i}$ and $\mathcal{E}^{h}:=\bigcup_{i=1}^{M_{e}}$ as the set of all edges in the coarse ad fine mesh $\mathcal{T}^{H}$ and $\mathcal{T}^{h}$, where $N_{e}$ and $M_{e}$ refer to the number of coarse/fine edges, respectively.
@@ Appendix A Knowledge of Mixed Generalized Multiscale Finite Element Methods for Darcy’s Equation[11]
[…]
use the space $\text{H}(\text{div},\Omega)=\{v=(v_{1},v_{2})\in(L^{2}(\Omega))^{2}:\ \nabla\cdot v\in L^{2}(\Omega)\}$ and define
@@
$\text{V}_{h}$ =$=\{\text{v}_{h}\in\text{V}:\ \text{v}_{h}|_{t}=(b_{t}x_{1}+a_{t},d_{t}x_{2}+c_{t}),\ a_{t},b_{t},c_{t},d_{t}\in\mathbb{R},t\in\mathcal{T}^{g}\}$
$\text{W}_{h}$ =$=\{\text{w}_{h}\in\text{W}:\text{$\text{w}_{h}$ is a constant on each element in $\mathcal{T}^{h}$}\}$
The normal components of $\text{v}_{h}$ are continuous across the interior edges in $\mathcal{T}^{h}$. Then the solution $(\text{u}_{h},\text{p}_{h})\in(\text{V}_{h},\text{W}_{h})$ on the fine grid will satisfy
$\int_{\Omega}k^{-1}\text{u}_{h}\cdot\text{v}_{h}-\int_{\Omega}\text{div}(\text{v}_{h})\text{p}_{h}$ $=0,\quad\quad\quad\ \ \forall\text{v}_{h}\in\text{V}_{h}^{0},$ (33)
$\int_{\Omega}\text{div}(\text{u}_{h})\text{w}_{h}$ $=\int_{\Omega}f\text{w}_{h},\quad\forall\text{w}_{h}\in\text{W}_{h}.$
where $\text{V}_{h}^{0}=\{\text{v}_{h}\in\text{V}_{h}:\text{v}_{h}\cdot n=0\ \text{on$\partial\Omega$}\}$. Then the system above can be written as the matrix form
$M_{\text{fine}}\text{U}_{h}+B_{\text{fine}}\text{P}_{h}$ $=0$ (34)
$B_{\text{fine}}^{T}\text{U}_{h}$ $=\text{F}_{h}$
@@
    Pressure Snapshot Space on the Find Grid: $\text{W}^{\text{snap}}=\{\phi_{i}^{\text{snap}}:\text{piecewise constant on the fine grid}\}$
@@
    $\begin{cases}k^{-1}\text{u}_{j}^{(i)}+\nabla\text{p}_{j}^{(i)}=0,\quad\text{in}\ T_{i},\\ \text{div}(\text{u}_{j}^{(i)})=0,\quad\quad\quad\ \ \ \text{in}\ T_{i},\\ \text{p}_{j}^{(i)}=\delta_{j}^{(i)}=\begin{cases}1\ \text{in}\ e_{j},\\ 0\ \text{on other fine-grid edges on}\ \partial\Omega\end{cases},j=1,2,\cdots J_{i}.\end{cases}$ (35)
@@
    $\begin{cases}k^{-1}\text{u}_{j}^{(i)}+\nabla\text{p}_{j}^{{(i)}}&=0,\ \ \text{in}\ T_{i},\\ \nabla\cdot\text{u}_{j}^{(i)}&=\alpha_{j},\ \text{in}\ T_{i},\\ \frac{\partial\text{p}_{j}^{(i)}}{\partial\text{n}_{i}}&=\delta_{j},\text{on}\ \partial\Omega\end{cases}$ (36)
    where $\text{n}_{i}$ refers to an outward unit norm vector on $\partial T_{i}$, $\alpha_{j}^{(i)}$ is chosen s.t. the compatibility condition $\int_{T_{i}}\alpha_{j}^{(i)}=\int_{\partial T_{i}}\delta_{j}^{(i)}$ is satisfied.
@@
$a_{i}(\text{p},\text{w})=\lambda s_{i}(\text{p},\text{w}),\ \forall\text{w}\in\text{W}^{\text{snap}}$ (37)
where $a_{i}(\text{p},\text{w})=\int_{T_{i}}k|\nabla\text{p}|^{2}$ is the stiffness matrix representation of snapshot space, $s_{i}(\text{p},\text{w})=\int_{T_{i}}k\text{pw}$ is the mass matrix representation. The discretization form of spectral problem can be written as:
@@
$\int_{\Omega}k^{-1}\text{u}_{\text{H}}\cdot\text{v}_{\text{H}}-\int_{\Omega}(\nabla\cdot\text{v}_{\text{v}_{\text{H}}})\text{p}_{\text{H}}$ $=0,\ \forall\text{v}_{\text{H}}\in\text{V}_{h},$ (40)
$\int_{\Omega}(\nabla\cdot\text{u}_{\text{H}})\text{w}_{\text{H}}$ $=\int_{\Omega}f\text{w}_{\text{H}},\ \forall\text{w}_{\text{H}}\in\text{W}^{{\text{off}}}$
After solving the coarse-grid pressure field $\text{p}_{\text{H}}$, the fine -grid pressure field $\text{p}_{\text{h}}$ will be obtained by interpolation using the offline basis functions:
$\text{p}_{h}=\text{R}_{\text{off}}\text{P}_{\text{H}}$ (41)
@@
Appendix B Examples of Reconstruction

### v2

@@ 4.2 Hierarchical Operator to Capture Global Information
Although the aforementioned model, by embedding physical laws and residual connections into a deep learning framework, can effectively approximate the solutions to PDEs, its primary convolutional structure, which focuses on local features, limits its ability to capture global interactions and long-range dependencies. To address this limitation, we have integrated a Transformer module. The Transformer enhances the model’s ability to capture global features and long-range dependencies through self-attention mechanisms and positional encoding, thereby improving the accuracy of predicting the pressure field $p_{\theta}(\cdot)$.
@@
$p_{\theta}(x)\xrightarrow{\scriptstyle\text{(Flatten)}}z\xrightarrow{\scriptstyle\text{(Pos Enc)}}z^{\prime}\xrightarrow{\scriptstyle\text{(Transformer)}}z^{\prime\prime}\xrightarrow{\scriptstyle\text{(FC)}}\hat{p}_{\theta}(x)$
@@
$\text{Multihead}(Q,K,V)=\text{Concat}(\text{head}_{1},\text{head}_{2},...,\text{head}_{h})W^{O}$ (25)
each $\text{head}_{i}$ is computed as $\text{Attention}(Q_{i},K_{i},V_{i})$, processing different linear projection subspaces, and $W^{O}$ is the output projection matrix belonging to $\mathbb{R}^{h\cdot d_{k}\times d}$, where $h$ is the number of attention heads, $d_{k}$ is the dimension of the key and query vectors, and $d$ is the dimension of the output.
@@
$\text{PE}(i,2k)=\sin\left(\frac{i}{10000^{\frac{2k}{d}}}\right),\quad\text{PE}(i,2k+1)=\cos\left(\frac{i}{10000^{\frac{2k}{d}}}\right)$
@@
In the context of solving Darcy Flow problem, this capability is particularly important. Certain parts of the pressure field in Darcy Flow may be influenced by the distribution of permeability and source terms at a distance. Therefore, the architecture of the Transformer model plays a significant positive role in understanding and predicting these complex spatial relationships.
@@ 4.3 Evaluation Metircs for this stage
Since we add physical information to the loss function, it is not appropriate to use the mean square error as an evaluation index directly. To do this, we use the $\text{R}^{2}$ and define the average absolute physical residual (AAPR) to evaluate the coefficient of variation of the model and the degree of fit to the data and physical information, which has been demonstrated by Equation.9.
@@
In this section, we propose a neural operator that integrates a convolutional operator with physical residual links and a Transformer module, designed to predict the pressure field from multi-scale basis functions and permeability fields. We define $\mathcal{G}_{\text{PINN}}$ as the first part of the mapping in Stage 2, mathematically expressed as:
@@
$\hat{p}_{\theta}(x)=\mathcal{G}_{\text{Total}}(X)=\mathcal{G}_{\text{Trans}}\left(\mathcal{G}_{\text{PINN}}(x)\right)=\int_{\Omega}\mathcal{K}\left(x,x^{\prime}\right)p_{\theta}\left(x^{\prime}\right)dx^{\prime}$ (31)
5 Numerical Results
@@
Table 1: Evaluation metrics for different basis functions.
MSE $\text{R}^{2}$
@@
As shown in the data presented in the table, the MSE of our proposed first-stage model on the test set for different basis functions is 0.0036, 0.0007, 0.0021, and 0.0027, respectively, while the corresponding R2 values are 0.9716, 0.9148, 0.9366, and 0.9542. These values indicate that such performance is acceptable for deep learning models. To further enhance the credibility of our results, we selected samples from the dataset used in this stage, specifically choosing coarse grids that include fractures and those without fractures, and reconstructed the corresponding basis functions for visualization. These results can be observed in the following figures: validation set—Fig.5 and 6; testing set—Fig.7 and 8; training set—Fig.12 and 14. It should be noted, however, that deep learning techniques cannot guarantee perfect learning of features from the data. Therefore, there is a possibility that the selected samples for visualization may include coarse grids with suboptimal reconstruction quality.
Table 2: R2 Performance of Different Basis Function Combinations.
Basis Functions Training Validation Testing Total
@@
For the second stage, reconstruction for the pressure field, the same step to illustrate the results. Tab.2 and 3 present the numerical results of the evaluation metrics used in our study. In this experiment, we combined various basis functions (including Basis 1, which requires no training) with their corresponding permeability fields and input them into the model for training. For our experiments, we randomly selected different numbers and combinations of basis functions for training. From the test set column in Tab.2, it can be observed that the R2 values of our model are all greater than 0.9, indicating that the model performs excellently in fitting the pressure data. Furthermore, as shown in Tab.3, the AAPR values are all at the magnitude of $1\times 10^{-4}$ or smaller, strongly demonstrating that the proposed model not only accurately reconstructs the data but also adheres to the underlying physical laws of real-world phenomena.
Table 3: AAPR Performance of Different Basis Function Combinations.
Basis Functions Training Validation Testing Total
@@
To comprehensively demonstrate the reconstruction capabilities of our method and the output characteristics of several key modules, we combined and presented these components. The corresponding results can be observed in Fig.10, 11 and 13. These visualizations include the permeability fields of high-contrast fractured porous media, the pressure field solutions of the Darcy equation, partial outputs of the preconditioner, and the reconstructed multiscale basis functions on the corresponding grids. This part of the results can fully show some key outputs of our proposed model and reflect the effectiveness of our model to a certain extent.
6 Discussion and Conclusion
In this paper, we proposed a hybrid-model-training two-stage method, developing an innovative neural operator to solve Darcy problem with high-contrast coefficients under the skeleton of mixed GMsFEM. In the first stage, traditional data-driven training method was hired to reconstruct the multiscale basis functions. For the purpose to eliminate the noise and expand the size of input features, we developed a domain transform-based preconditioner, using FNO to process the high-contrast permeability fields, together with a hierarchical multiscale feature extractor to capture the information contained. In the second stage, to efficiently solving the pressure field, we used physics-informed training method to achieve our purpose with a small-sample dataset. To utilize the reconstructed basis functions, we concatenate them with permeability fields as the model input. They will be processed by both local (symmetrical residual connections-based accelerator with a V-shaped CNN) and global (Transformer-based for capture spatial information and long-range dependencies) information learner. Both stages achieved good results in the reconstruction of multiscale basis functions and pressure fields.
Compared to existing research, our proposed model demonstrates superior performance. For the reconstruction of basis functions, since our model is primarily trained in a data-driven manner, the evaluation metrics focusing on the data itself are of particular importance. On the test set, our model achieved mean squared error (MSE) values of 0.0036, 0.0007, 0.0021, and 0.0027, and corresponding R2 values of 0.9716, 0.9148, 0.9366, and 0.9542 for different basis functions. In contrast, the results from [14] reported MSE values of 0.0466, 0.0743, 0.0184, and 0.0154, with corresponding R2 values of 0.8083, 0.6445, 0.8762, and 0.7625. These results indicate that our model achieves better data fitting and exhibits stronger generalization capabilities.
For the reconstruction of pressure fields in the second stage, our model achieved R2 values ranging from 0.9032 to 0.9111 for different basis function combinations. In comparison, [15] reported R2 values ranging from 0.8456 to 0.9191. While the R2 values are comparable, we place greater emphasis on whether the results align with real-world physical laws. To address this, we incorporated the average absolute percentage residual (AAPR) as an additional evaluation metric. In this context, all combinations exhibited AAPR values on the order of $1\times 10^{-4}$ or smaller. Despite the similar R2values, our model achieves a higher lower bound for R2, indicating that it delivers comparable or better reconstruction performance while better adhering to physical laws.
@@ 6 Discussion and Conclusion
While the proposed framework performs well in reconstructing multiscale basis functions and pressure fields, several limitations should be noted. The reliance on large, high-quality datasets in the first stage may limit its applicability in scenarios where data is scarce or expensive to acquire. In addition, the inclusion of physical constraints in the second stage, while beneficial for maintaining physical consistency, adds complexity to the training process and requires careful tuning of hyperparameters. The computational overhead introduced by the Transformer module also presents challenges when scaling to larger datasets or higher dimensional domains.
@@
Appendix A Knowledge of Mixed Generalized Multiscale Finite Element Methods for Darcy Equation
We consider the following first-order Darcy’s Flow problem with high-contrast coefficients and nonhomogeneous boundary condition in Lipschitz continuous domain:
$\begin{cases}\kappa^{-1}\mathbf{u}+\nabla p&=0,\quad\text{in}\,\Omega\\ \nabla\cdot\mathbf{u}&=f,\quad\text{in}\,\Omega\\ \mathbf{u}\cdot\mathbf{n}&=g,\quad\text{on}\,\partial\Omega\end{cases}$ (32)
where $\kappa$ is high-contrast heterogeneous permeability field, $\mathbf{u}$ and $p$ are the Darcy velocity and pressure, respectively. $f$ is the known source term, $g$ is given normal component of Darcy velocity on the boundary, $\Omega=[0,1]^{2}$ is the computational domain, and $\mathbf{n}$ is the outward unit norm vector on the boundary.
Let $\mathcal{T}^{H}$ denote a conforming partition of $\Omega$ into finite elements with coarse-grid size $H$ and $\mathcal{T}^{h}$ denote the fine-grid partition of $\Omega$ into non-overlapping elements with size $h$. Define $\mathcal{E}^{H}:=\bigcup_{i=1}^{N_{e}}E_{i}$ and $\mathcal{E}^{h}:=\bigcup_{i=1}^{M_{e}}e_{i}$ as the set of all edges in the coarse ad fine mesh $\mathcal{T}^{H}$ and $\mathcal{T}^{h}$, where $N_{e}$ and $M_{e}$ refer to the number of coarse and fine edges, respectively.
@@ Appendix A Knowledge of Mixed Generalized Multiscale Finite Element Methods for Darcy Equation
$L^{2}(\Omega)=\{v:\text{$v$ is defined on $\Omega$~\text{and}~$\int_{\Omega}v^{2}\text{d}x<\infty$ \}}$
Use the space $\text{H}(\text{div},\Omega)=\{\mathbf{v}=(\mathbf{v}_{1},\mathbf{v}_{2})\in(L^{2}(\Omega))^{2}:\ \nabla\cdot\mathbf{v}\in L^{2}(\Omega)\}$. Define
@@
$\text{V}_{h}$ $=\{\mathbf{v}_{h}\in\text{V}:\ \mathbf{v}_{h}|_{t}=(b_{t}x_{1}+a_{t},d_{t}x_{2}+c_{t}),\ a_{t},b_{t},c_{t},d_{t}\in\mathbb{R},t\in\mathcal{T}^{g}\}$
[…]
The normal components of $\mathbf{v}_{h}$ are continuous across the interior edges in $\mathcal{T}^{h}$. Then the solution $(\mathbf{u}_{h},p_{h})\in(\text{V}_{h},\text{W}_{h})$ on the fine grid will satisfy
$\int_{\Omega}\kappa^{-1}\mathbf{u}_{h}\cdot\mathbf{v}_{h}-\int_{\Omega}\text{div}(\mathbf{v}_{h})p_{h}$ $=0,\quad\quad\quad\ \ \forall\mathbf{v}_{h}\in\text{V}_{h}^{0},$ (33)
$\int_{\Omega}\text{div}(\mathbf{u}_{h})\text{w}_{h}$ $=\int_{\Omega}f\text{w}_{h},\quad\forall\text{w}_{h}\in\text{W}_{h}$
where $\text{V}_{h}^{0}=\{\mathbf{v}_{h}\in\text{V}_{h}:\mathbf{v}_{h}\cdot n=0\ \text{on $\partial\Omega$}\}$. Then the system above can be written as the matrix form
$\text{M}_{\text{fine}}\text{U}_{h}+\text{B}_{\text{fine}}\text{P}_{h}$ $=0$ (34)
$\text{B}_{\text{fine}}^{T}\text{U}_{h}$ $=\text{F}_{h}$
@@
    Pressure Snapshot Space on the Find Grid:
    $\text{W}^{\text{snap}}=\{\phi_{i}^{\text{snap}}:\text{piecewise constant on the fine grid}\}$
@@
    $\begin{cases}\kappa^{-1}\mathbf{u}_{j}^{(i)}+\nabla p_{j}^{(i)}=0,\quad\text{in}\ T_{i},\\ \text{div}(\mathbf{u}_{j}^{(i)})=0,\quad\quad\quad\ \ \ \text{in}\ T_{i},\\ p_{j}^{(i)}=\delta_{j}^{(i)}=\begin{cases}1\ \text{in}\ e_{j},\\ 0\ \text{on other fine-grid edges on}\ \partial\Omega\end{cases},j=1,2,\cdots J_{i}.\end{cases}$ (35)
@@
    $\begin{cases}\kappa^{-1}\mathbf{u}_{j}^{(i)}+\nabla p_{j}^{{(i)}}&=0,\ \ \text{in}\ T_{i},\\ \nabla\cdot\mathbf{u}_{j}^{(i)}&=\alpha_{j},\ \text{in}\ T_{i},\\ \frac{\partial p_{j}^{(i)}}{\partial\mathbf{n}_{i}}&=\delta_{j},\text{on}\ \partial\Omega\end{cases}$ (36)
    where $\mathbf{n}_{i}$ refers to an outward unit norm vector on $\partial T_{i}$, $\alpha_{j}^{(i)}$ is chosen s.t. the compatibility condition $\int_{T_{i}}\alpha_{j}^{(i)}=\int_{\partial T_{i}}\delta_{j}^{(i)}$ is satisfied.
@@
$a_{i}(p,\text{w})=\lambda s_{i}(p,\text{w}),\ \forall\text{w}\in\text{W}^{\text{snap}}$ (37)
where $a_{i}(p,\text{w})=\sum_{e\in\mathcal{E}_{h}^{0}}\kappa_{e}[p_{h}]_{e}[\text{w}_{h}]_{e}$ is the stiffness matrix representation of snapshot space, $s_{i}(p,\text{w})=\int_{T_{i}}kp\text{w}$ is the mass matrix representation. The discretization form of spectral problem can be written as:
@@
$\int_{\Omega}\kappa^{-1}\mathbf{u}_{\text{H}}\cdot\mathbf{v}_{\text{H}}-\int_{\Omega}(\nabla\cdot\mathbf{v}_{{\text{H}}})p_{\text{H}}$ $=0,\ \forall\mathbf{v}_{\text{H}}\in\text{V}_{h},$ (40)
$\int_{\Omega}(\nabla\cdot\mathbf{u}_{\text{H}})\text{w}_{\text{H}}$ $=\int_{\Omega}f\text{w}_{\text{H}},\ \forall\text{w}_{\text{H}}\in\text{W}^{{\text{off}}}$
After solving the coarse-grid pressure field $p_{\text{H}}$, the fine-grid pressure field $p_{\text{h}}$ will be obtained by interpolation using the offline basis functions:
$p_{h}=\text{R}_{\text{off}}\text{P}_{\text{H}}$ (41)
@@
Appendix B Examples of Our Computation Results